# Human-free Prompted Based Anomaly Detection: prompt optimization with Meta-guiding prompt scheme

### v1

@@ Abstract
Pre-trained vision-language models (VLMs) are highly adaptable to various downstream tasks through few-shot learning, making prompt-based anomaly detection a promising approach. Traditional methods depend on human-crafted prompts that require prior knowledge of specific anomaly types. Our goal is to develop a human-free prompt-based anomaly detection framework that optimally learns prompts through data-driven methods, eliminating the need for human intervention. The primary challenge in this approach is the lack of anomalous samples during the training phase. Additionally, the Vision Transformer (ViT)-based image encoder in VLMs is not ideal for pixel-wise anomaly segmentation due to a locality feature mismatch between the original image and the output feature map. To tackle the first challenge, we have developed the Object-Attention Anomaly Generation Module (OAGM) to synthesize anomaly samples for training. Furthermore, our Meta-Guiding Prompt-Tuning Scheme (MPTS) iteratively adjusts the gradient-based optimization direction of learnable prompts to avoid overfitting to the synthesized anomalies. For the second challenge, we propose Locality-Aware Attention, which ensures that each local patch feature attends only to nearby patch features, preserving the locality features corresponding to their original locations. This framework allows for the optimal prompt embeddings by searching in the continuous latent space via backpropagation, free from human semantic constraints. Additionally, the modified locality-aware attention improves the precision of pixel-wise anomaly segmentation.
@@ 1 Introduction
In this paper, we propose a human-free prompting anomaly detection framework, ”Meta-prompting Semantic Learning for Anomaly Detection with Locality-aware Attention Image Encoder”. The proposed framework follows the intuition of data-driven methods that leverage back-propagation to actively find the prompt that can generate the optimal textual embedding suited for anomaly detection, instead of passively combining all possible human semantic prompts to meet the requirements of prompt-based anomaly detection. To address the limitation of the absence of anomaly samples for context optimization, we develop the Object-Attention Anomaly Generation Module to synthesize anomaly samples for training. Meanwhile, the proposed Meta-guiding Prompt-tuning Scheme (MPTS) is developed to prevent the learnable prompt from overfitting on synthesized anomalies through iterative gradient-based calibration.
@@
Additionally, we introduce Locality-aware Attention, which can be adopted by any transformer-based model to mitigate the misalignment of input token features and corresponding output token features. This mechanism restricts attention to local neighborhoods, preserving essential local details and enhancing the alignment between the input features and the output tokens. By focusing on neighboring regions, the locality-aware attention ensures that the feature extraction process maintains the spatial integrity of the image. This addresses the limitations of the vanilla attention mechanism, where distant tokens tend to dilute crucial local information, resulting in the misalignment of locality features between input and output. aThe contributions are listed as follows:
  • We propose a novel paradigm for a prompt-based anomaly detection framework, namely Meta-prompting Semantic Learning for Anomaly Detection with Locality Feature Attention Image Encoder, which optimizes the pre-trained VLM for anomaly detection tasks in a human-free prompting manner.
  • We introduce the Meta-guiding Prompt-tuning Scheme along with the Object-Attention Anomaly Generation Module to enable the optimization of learnable prompts in the absence of anomaly samples. The optimized learnable prompts with the Meta-guiding Prompt-tuning Scheme even exceed manually designed prompts by a significant margin.
  • We propose the Locality-Aware Transformer, which aims to extract a feature map where all locality features align with their original image positions, ensuring pixel-wise anomaly segmentation in the CLIP-based framework.

### v4

# Human-Free Automated Prompting for Vision-Language Anomaly Detection: Prompt Optimization with Meta-guiding Prompt Scheme

@@ Abstract
Pre-trained vision-language models (VLMs) are highly adaptable to various downstream tasks through few-shot learning, making prompt-based anomaly detection a promising approach. Traditional methods depend on human-crafted prompts that require prior knowledge of specific anomaly types. Our goal is to develop a human-free prompt-based anomaly detection framework that optimally learns prompts through data-driven manner, eliminating the need for human intervention. The primary challenge in this approach is the lack of anomalous samples during the training phase. To tackle this challenge, we develop the Object-Attention Anomaly Generation Module (OAGM) to synthesize anomaly samples for training. Furthermore, to prevent learned prompt from overfit on synthesized anomaly feature, we proposed Meta-Guiding Prompt-Tuning Scheme (MPTS) that iteratively adjusts the gradient-based optimization direction of learnable prompts to avoid overfitting to the synthesized anomalies . This framework allows for the optimal prompt embeddings by searching in the continuous latent space via backpropagation, free from human semantic constraints. Additionally, the modified locality-aware attention improves the precision of pixel-wise anomaly segmentation.
@@ 1 Introduction
In this paper, we propose a human-free automated prompting anomaly detection framework, namely ”Meta-prompting Semantic Learning for Anomaly Detection with Locality-aware Attention Image Encoder”. The proposed framework follows the intuition of data-driven methods that leverage back-propagation to actively find the prompt that can generate the optimal textual embedding suited for anomaly detection, instead of passively combining all possible human semantic prompts to meet the requirements of prompt-based anomaly detection. To address the limitation of the absence of anomaly samples for context optimization, we develop the Object-Attention Anomaly Generation Module to synthesize anomaly samples for training. Meanwhile, the proposed Meta-guiding Prompt-tuning Scheme (MPTS) is developed to prevent the learnable prompt from overfitting on synthesized anomalies through iterative gradient-based calibration.
@@
Additionally, we introduce Locality-aware Attention, which can be adopted by any transformer-based model to mitigate the misalignment of input token features and corresponding output token features. This mechanism restricts attention to local neighborhoods, preserving essential local details and enhancing the alignment between the input features and the output tokens. By focusing on neighboring regions, the locality-aware attention ensures that the feature extraction process maintains the spatial integrity of the image. This addresses the limitations of the vanilla attention mechanism, where distant tokens tend to dilute crucial local information, resulting in the misalignment of locality features between input and output. The contributions are listed as follows: (1) We propose a novel paradigm for a prompt-based anomaly detection framework, namely Meta-prompting Semantic Learning for Anomaly Detection with Locality Feature Attention Image Encoder, which optimizes the pre-trained VLM for anomaly detection tasks in a human-free prompting manner; (2) We introduce the Meta-guiding Prompt-tuning Scheme along with the Object-Attention Anomaly Generation Module to enable the optimization of learnable prompts in the absence of anomaly samples. The optimized learnable prompts with the Meta-guiding Prompt-tuning Scheme even exceed manually designed prompts by a significant margin; (3) We propose the Locality-Aware Transformer, which aims to extract a feature map where all locality features align with their original image positions, ensuring pixel-wise anomaly segmentation in the CLIP-based framework.